OpenStudy (anonymous):

x^2+5x=36 Solve each quadratic equation using factoring?? Pleasee help :)

OpenStudy (anonymous):

(x+9)(x-4)=0 and then set factors = to 0 to solve

OpenStudy (anonymous):

@lisapeajac you just wanna help me with all of them? There's only like 9 left :)

OpenStudy (anonymous):

Sure. Factoring is solving a puzzle. What multiplies to give you last number and adds/subtracts to give you the middle number

OpenStudy (anonymous):

okay.. umm @lisapeajac 2x^2+9x-5=0

OpenStudy (theviper):

\(\Large{x^2+5x-36=0}\) \(\Large{=>x^2+9x-4x-36=0}\) \(\Large{=>x(x+9)-4(x+9)=0}\) \(\Large{=>(x-4)(x+9)}\)

OpenStudy (anonymous):

thanks @TheViper

OpenStudy (theviper):

You r welcome ;)